OpenStudy (anonymous):

Find an equation of variation where y varies directly as x and y = 2 when x = 10, find y when x = -25.

jimthompson5910 (jim_thompson5910):

y varies directly as x, so y = kx for some constant k They give you y = 2 when x = 10, which means 2 = k*10 From here, solve for k. Let me know what you get.

OpenStudy (anonymous):

y=5

jimthompson5910 (jim_thompson5910):

no, and it's not y = 5 as it should be k = something

jimthompson5910 (jim_thompson5910):

2 = k*10 2/10 = k 1/5 = k k = 1/5

OpenStudy (anonymous):

on my paper it has y= allready there

jimthompson5910 (jim_thompson5910):

well finding k is an intermediate step

jimthompson5910 (jim_thompson5910):

you don't report the value of k in this case

jimthompson5910 (jim_thompson5910):

so because y = kx and k = 1/5, we know that y = kx y = (1/5)x From here, plug in x = -25 and evaluate to find y

OpenStudy (anonymous):

anyone help me

jimthompson5910 (jim_thompson5910):

what do you get when you plug x = -25 into y = (1/5)x

OpenStudy (anonymous):

-5

jimthompson5910 (jim_thompson5910):

good you got it

OpenStudy (anonymous):

Thanks @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome